# Towards Social AI: A Survey on Understanding Social Interactions

### v1

@@ IV-C5 Discussion
V Benchmark Datasets
V-A Datasets with Verbal Cues
V-A1 Dialogue Act Analysis Datasets
Different datasets on dialogue act analysis usually focus on different domains such as train booking, restaurant search, counseling and so on. They are labeled with domain-specific dialogue acts, and systems are evaluated using the accuracy in correctly classifying the dialogue act labels.
SwitchboardDA Switchboard Dialog Act Corpus [22] contains 1,155 five-minute dialogues between two speakers on telephone, resulting in a total of 221,616 utterances. It contains speaker-level, utterance-level and conversation-level of information and is annotated with approximately 60 different speech act tags.
MRDA Corpus ICSI Meeting Recorder Dialog Act Corpus [261] consists of 72 hours of speech with over 180,000 manually annotated dialog act tags. The dialogues happen in meetings with an average of six speakers. The corpus is annotated with dialogue act segment boundaries, dialogue acts, and adjacency pairs.
AMI Corpus [262] contains 100 hours of video recordings of scenario-driven meetings. Participants were assigned the roles of employees in a company who were discussing the development of a new product. The dialogues consist of four phases: project kick-off, functional design, conceptual design, and detailed design. The data comes with annotations in terms of transcriptions, dialogue acts and summaries.
DSTC Dialog State Tracking Challenge dataset [263] provides a collection of 1,500 human-computer dialogues with a set of 11 evaluation metrics. It uses the conversation data from the public deployment of several systems. Each dialogue is annotated with dialogue states that represent user’s goals and intentions.
DSTC2 The second Dialog State Tracking Challenge [264] introduces some new features. It introduces a new domain of restaurant search. The users’ goals are also allowed to change during the dialogues. Moreover, it uses a richer representation of dialog states which covers the search method and information the user would like the system to read out. It contains 3235 dialogues in total.
DSTC3 The third Dialog State Tracking Challenge [265] contains 2,275 dialogues from paid crowdworkers. It introduces a new domain on tourist information. Also, there are new entity types and values in the test data which do not appear in the train data, posing greater challenges in terms of generalization capability of models.
HOPE [20] contains around 12,900 utterances from counseling transcripts from publicly available Youtube videos. The texts are annotated with 12 different dialogue act labels. They can be grouped into speaker initiative labels, speaker responsive labels and general labels.
IMCS-21 [266] is a dialogue dataset on medical consultation consisting of 4,116 annotated dialogues with 164,731 utterances. The data comes from online professional medical consulting community. There are two broad dialogue act categories, which are inform and request. It is also annotated with more fine-grained categories such as basic information, symptom, etiology, existing exam and treatment.
V-A2 Dialogue Emotion Analysis Datasets
Researchers have created various dialogue datasets for emotion analysis. The dialogues can be dyadic and multi-party. The number of emotion classes in each dataset can vary depending on different annotation schemes. Systems are evaluated using the accuracy score and F1 score in classifying the emotions.
DailyDialogue [267] consists of 13,118 human-written high-quality dialogues. The dialogues are in the multi-turn format with rich contextual information, where most previous dialogue datasets did not have such format. The dialogues are daily conversations on a variety of topics across multiple domains. They are annotated with communication functions and six categories of emotions.
EmotionLines [42] contains 2000 multi-party dialogues from Friends TV scripts, resulting in a total of 29,245 utterances. They are labeled with six Ekman’s basic emotions and one neutral emotion by five Amazon MTurkers.
EmoContext [268] was published as a task in SemEval-2019. It presents the dialogue context with two previous turns and contains annotation of four emotion classes (Happy, Sad, Angry and Others). There is a total of 38,424 dialogues.
EmoWOZ [269] is built based on previous established multi-domain task-oriented dialogue dataset MultiWOZ [276]. On top of MultiWOZ, the authors collected human-machine dialogues in the same domains to ensure a better coverage of different emotion types. In total, there are more than 11,000 dialogues and 83,000 emotion annotations.
V-A3 Common Sense Reasoning
Common sense reasoning benchmarks usually contain a context such as event and premise, provide possible options for disambiguation of coreference and possible inference on people’s mental states and plausible following events. They evaluate systems’ inference capability using classification accuracy in terms of accuracy scores and F1 scores.
WS Challenge Winograd Schema Challenge [270] is built by selecting one sentence from each sentence pair created under Winograd schema. Sentences in the same pair are different from each other in just one or two words, which lead to different referential resolutions in the sentences. Models need common sense reasoning skills to achieve human-level accuracy in terms of resolving the ambiguities correctly.
SNLI Stanford Natural Language Inference [271] holds a collection of 570,000 sentence pairs consisting of a premise and a hypothesis, which are annotated with the logical connection in terms of entailment, contradiction and neutral. Each pair is annotated by five different crowdworkers. Understanding the semantic relationship between two sentences in a pair require the common sense reasoning ability.
Story Cloze Test [272] contains stories randomly sampled from ROCStories Corpus. The crowdworkers will be presented with the first four sentences of each story and annotate the story with a right ending and a wrong ending based on common sense. The annotated stories will be verified by three crowdworkers for quality control, resulting in a final set of 3,742 test cases.
JOCI JHU Ordinal Commonsense Inference [273] contains 39,093 sentence pairs annotated with an ordinal scale of plausibility. The authors sampled contexts from SNLI and ROCStories and generated possible inference candidates for those contexts using general world knowledge or neural methods. After that, the sentence pairs were annotated with level of plausibility via crowdsourcing.
Story Commonsense [274] holds a collection of 15,000 annotated stories under a carefully designed annotation pipeline. The stories have annotations on entities, actions, affects, motivations and emotion reactions. It can benchmark models’ common sense reasoning performance in classifying motivations and emotion reactions and providing explanations.
Event2Mind [78] consists of 24,716 phrases on a wide variety of daily events. The events are extracted from ROCStories, Google Syntactic N-grams and Spinn3r corpus. They are annotated with possible intents and reactions of the events’ participants via crowdsourcing.
CommonsenseQA [275] contains 12,247 questions that require the common sense of real world knowledge. The dataset contains data in MCQ formats. The instances are designed in a challenging way where similar concepts in ConceptNet are extracted and semantically difficult questions are designed by crowdworkers.
SocialIQA [60] is on common sense reasoning for social situations. There are 38,000 MCQs in the dataset about social situations. The questions and answers are collected from crowdworkers. To address the potential problem of annotation artifacts during crowdsourcing, the authors combined manually created negative answers and adversarial question-switched answers.
@@ V-A3 Common Sense Reasoning
V-B Datasets with Non-verbal Cues
V-B1 Gesture Datasets
The majority of gesture recognition datasets are based on video data. Recent advancements in research underscore a new trend by increasing the distance between people and the capturing devices to increase models’ robustness [109, 282]. In this subsection, we introduce 9 public benchmarks and datasets crucial for gesture recognition research. The evaluation metrics used to assess model efficacy on these benchmarks include some classification metrics like Accuracy, Mean Average Precision (mAP), and the F1-score. For tasks related to continuous gesture recognition [277, 109], the Jaccard Index or Mean Jaccard Index are employed to provide a thorough evaluation of a model’s ability to accurately segment continuous gestures. The Jaccard Index measures the proportion of overlap between the actual gesture segment and the model’s predicted segment over the total number of frames, providing a quantitative measure of accuracy for gesture recognition.
ChaLearn IsoGD Both ChaLearn IsoGD and ChaLearn ConGD [277] trace their origins to the CGD 2011 [292]. The ChaLearn IsoGD dataset is a large-scale dataset dedicated to isolated hand gestures. It encompasses 47,933 RGB-D gesture videos. Each RGB-D video uniquely represents a single gesture, with a total of 249 gesture labels performed by 21 different individuals. Each gesture category is represented by over 200 RGB and depth videos.
ChaLearn ConGD [277], developed alongside the ChaLearn IsoGD dataset, focuses on continuous gesture recognition. It comprises 22,535 RGB-D videos, capturing a total of 47,933 gestures. Unlike isolated gestures, each RGB-D video in the ChaLearn ConGD dataset may encompass one or several gestures, spanning across 249 distinct gesture labels executed by 21 different participants.
EgoGesture [278] is a large-scale, multimodal collection designed for egocentric perspective. It encompasses 2,081 RGB-D videos from 50 distinct subjects, containing 24,161 gesture instances and 2,953,224 frames. The dataset features 83 types of static or dynamic gestures, tailored for interactions with wearable devices. These videos were collected from 6 varied indoor and outdoor settings.
Jester [279] features 148,092 video clips annotated with 27 gesture categories, showcasing individuals executing predefined gestures in front of web cameras, including actions like two-finger swipes down, left or right swipes, and finger taps. The inclusion of two “no gesture” categories enhances the model’s ability to differentiate between specific gestures and incidental hand movements.
IPN Hand [280] comprises over 4,000 instances of hand gestures across 800,000 frames from 50 participants. It encompasses 13 types of static and dynamic gestures designed for interactions with non-touchscreen interfaces. The IPN Hand features the highest rate of intra-class variation and also offers real-time optical flow and hand segmentation results.
iMiGUE [82] is the first public dataset for emotional micro-gestures, which are involuntary actions motivated by underlying emotions. The dataset is annotated on two levels: the micro-gesture categories were annotated on video clip-level, and the emotion categories were labeled on video-level. The iMiGUE organizes 32 distinct categories for 359 videos from post-match press conferences.
LD-ConGR [109] features a total of 542 videos and 44,887 instances of hand gestures. These videos were collected from 30 participants across five distinct environments, using a Kinect V4 camera to capture footage from a third-person perspective. Each video comprises both a color stream and a depth stream, which are synchronized.
SMG [281], designed for micro-gesture recognition, consists of 40 video sequences (8 hours total) from 40 participants. It includes RGB, depth, silhouette videos, and skeleton data. The dataset contains 3712 pre-segmented and labeled micro-gesture clips across 16 classes and non-micro-gesture classes. It also includes 71 instances of relaxed and stressed emotional states, each lasting 2-5 minutes.
HaGRID [282] is a large-scale, high-resolution dataset for static gesture classification or detection. It contains 552,992 FullHD RGB images across 18 hand gesture classes. The dataset was primarily collected indoors, with some images taken under extreme lighting conditions. Participants performed gestures at distances ranging from 0.5 to 4 meters from the camera for diverse spatial representations.
V-B2 Gaze Datasets
Various gaze analysis datasets address the tasks of predicting spatial gaze targets, where predictions are evaluated against ground truth gaze points, and recognizing multi-person gaze behaviors, where predictions are evaluated at the frame or instance level.
GazeFollow [110] consists of 122,143 internet images annotated with head bounding boxes and gaze points. The train set has one gaze annotation per instance, while the test set has $\approx 10$ unique annotations per instance. In total, there are 130,339 instances (head bounding box with paired gaze target).
VideoGaze [126] addresses the unique case of identifying gaze targets in following video frames, where the camera often pans from the person to the gaze target. Each annotation consists of 6 frames, where the head bounding box is provided for the first frame and the gaze target locations are annotated for the following 5 frames (including indication of if the target is out of frame). It is sourced from 140 movies.
DLGaze [120] consists of 86 videos manually collected in 4 unique locations by 16 people. The participants themselves then annotated the location of their gaze targets. In total there are 95,000 frames. 4 frames are annotated per second, so there are $\approx 6,333$ instances.
VideoCoAttention [151] is a video dataset for identifying shared attention, which occurs when two or more people are looking at the same target. Shared attention instances are annotated as a set of head bounding boxes for the people involved and a bounding box for the shared target of gaze. In total, there are 138,203 instances of shared attention annotated across 492,1000 frames from 380 videos of TV shows / movies.
EGTEA Gaze+ [283] is an egocentric dataset consisting of 28 hours of people cooking in a kitchen environment. It is collecting with wearable eye tracking glasses, providing ground truth for egocentric gaze estimation.
VACATION [165] is a video dataset for recognizing a broader set of multi-party communicative gaze behaviors. It defines 6 atomic-level behaviors: single, mutual, avert, refer, follow, and share, as well as 5 event-level gaze behaviors: non-communicative, mutual gaze, gaze eversion, gaze following, and joint attention. The dataset contains 300 video sequences from movies and TV shows, with a total of 96,993 frames.
UCO-LAEO & AVA-LAEO [160] are benchmarks for identifying mutual gaze, or “looking at each other”. They contain a total of 208,688 instances comprised of head bounding boxes for two individuals and a binary label indicating whether they are looking at each other. UCO-LAEO has $\approx 18,000$ frames from 4 TV shows, while AVA-LAEO includes $\approx 1.4M$ frames from 298 movies, sourced from the AVA dataset [293].
VideoAttentionTarget [111] is comprised of 606 video segments from 50 TV shows or movies. Annotations consist of head bounding box tracks with an annotated gaze point. Additionally, VideoAttentionTarget labels if the gaze point is in or out of frame and includes this as an additional benchmark task. In total, there are 71,666 frames and 164,541 instances.
Triadic Belief [294] contains 88 videos (109,331 frames) of social interactions captured from both third-person and first-person views. It’s annotated with gaze, gestures, communication events, and belief dynamics.
Ego4D [146] contains 31 hours of egocentric video collected with wearable eye tracking glasses, which is used as a benchmark for egocentric gaze estimation. This eye tracking subset was collected in multi-person social settings.
ChildPlay [112] targets on the specific domain of children interacting with adults and other children, sourcing videos from Youtube. Annotations follow the same scheme as VideoAttentionTarget, consisting of the head bounding box, annotated gaze point, and in vs. out of frame label.
V-B3 Facial Expression Datasets
Having sufficiently large and labelled dataset is essential to ensure the optimal learning of deep facial expression recognition (FER) methods. This dataset should encompass a diverse range of variations in both the surrounding environment and facial structures. In this section, we delve into the discussion of public benchmark FER databases, commonly utilized in the researched papers, which cover fundamental emotions, which are classified into seven categories: happy, anger, fear, disgust, sad, surprise, neutral. The performance is evaluated using several metrics, including precision, accuracy, recall, specificity, and F1-score.
MMI [178, 179] is a laboratory-controlled dataset containing 326 sequences from 32 subjects. It includes 213 sequences annotated with six basic expressions, with 205 sequences captured in a frontal view. MMI sequences start from a neutral expression, reach a peak, and return to a neutral state. The dataset presents challenges due to inter-personal variations, non-uniform expressions, and subjects wearing accessories.
CK+ [284] is a facial expression dataset captured in a controlled lab environment. It contains 593 video sequences of 123 subjects, with durations ranging from 10 to 60 frames, showing the transition from neutral to peak emotional expressions. 327 sequences from 118 subjects are annotated with seven basic emotion labels using the Facial Action Coding System (FACS).
FER-2013 [285] consists of 35,887 grayscale images categorized into seven different facial expressions. The images are relatively low-resolution (48x48 pixels) and depict diverse individuals across different age groups and ethnicities, consisting of both posed and unposed headshots.
SFEW The Static Facial Expressions in the Wild [286] dataset was generated by selecting static frames from the AFEW database through the computation of key frames based on facial point clustering. Primarily used as benchmark data for the SReco sub-challenge in EmotiW 2015, it consists of 1,766 samples. Each image is labeled with one of seven expression categories
EmotioNet EmotioNet [287] is an extensive dataset that comprises one million facial images sourced from the Web. 950,000 umages were annotated using an automatic action unit detection model whereas the remaining 25,000 images underwent manual annotation, specifically identifying 11 action units (AUs).
AFEW The Acted Facial Expressions in the Wild [288] contains video clips from movies with spontaneous expressions, diverse head poses, occlusions, and varying illuminations. It serves as a temporal and multimodal database with audio and video data. The dataset includes samples labeled with seven expressions from reality TV shows. For EmotiW 2017, AFEW 7.0 consists of 1,809 samples.
RAF-DB Real-world Affective Faces Database [289, 290] is a comprehensive dataset containing 29,672 facial images with basic or compound expression labels, annotated by 40 independent taggers. The dataset features diverse subjects, varied head poses, lighting conditions, occlusions, and post-processing effects, making it a rich resource for studying facial expression recognition in real-world scenarios.
AffectNet AffectNet [180] comprises over one million images sourced from Web by querying different search engines with keywords related to facial expressions. 450,000 images have manually annotated for eight expressions, and it is the largest dataset providing images with facial expressions for both categorical and dimensional emotional models.
ExpW The Expression in-the-Wild (ExpW) [200] dataset is a collection of facial images specifically labeled for expressions. The dataset includes 91,793 faces, each manually annotated with one of seven basic expression categories. Notably, ExpW stands out for its extensive size and diverse facial variations, surpassing many existing databases in terms of the breadth of facial expressions captured.
Face MPMI The Face Multi-Pose Multi-Illumination (Face MPMI) [291] dataset contains dynamic facial expressions from 104 subjects, captured under 13 pose variations, 4 illumination conditions (room, bright, left, and right), and with/without eyeglasses. This comprehensive dataset enables evaluation under diverse challenging conditions.

### v2

@@ 2.3.3 Discussion
2.4 Datasets with Verbal Cues
Table I shows a comprehensive overview, comparing the attributes of datasets with verbal social cues. Further details for each dataset are included in the supplementary material.
3 Understanding Non-verbal Cues
3.1 Gesture and Body Language Analysis
3.1.1 Social Background
When people engage in face-to-face communication, hand movements often accompany their words to emphasize points or release tension. These instinctive hand motions, synchronized with speech, are termed co-speech gestures, and they naturally enhance languages. Across different cultures and linguistic backgrounds, gestures and body language are integral to communication [93, 94]. Notably, babies gesture even before uttering their first words [95]. Theoretical models posit that both speech and gesture spring from a shared representational system [96, 97]. This perspective suggests that comprehensive conceptual frameworks encompass both imagery-driven and symbolic data, such as gestures and speech, respectively [98, 99]. McNeill et al. [100] identified four types of gesture: deictic (e.g., pointing), beat (e.g., rhythmic emphasis), iconic (e.g., mimicking form), and metaphoric (e.g., abstract).
3.1.2 Isolated Gesture Recognition
Early works in gesture recognition mainly addressed isolated gestures, which are standalone hand motions or signs. They are identified as separate entities, without considering a sequence of multiple gestures. Early works in isolated gesture recognition utilized optical flow and human pose analysis. The emergence of CNN and 3D CNN has made certain breakthroughs in this field [101, 102]. Lee et al. [103] utilized CNN extract feature maps from static images and then proposed a motion filter to merge 2-frame motion information to predict isolated gestures. Zhou et al. [104] further proposed the framework of temporal relation networks for utilizing temporal relations among multiple CNN-based video frames, extending the idea of relation network [105] into gesture recognition. Recent work [106] attempted to mitigate the impact of gesture-irrelevant information (e.g., background, illumination) by introducing information-theoretic loss for isolated gesture recognition. Research on 3D skeleton gesture recognition focusing on isolated gestures has also gained considerable attention, as demonstrated by the works of De Smedt et al. [107], who utilized Fisher vector representation with Gaussian Mixture Models. Additionally, Aich et al. [108] have proposed a strategy by developing a boundary-aware prototypical sampling technique. This method improves model inversion for class-incremental learning in 3D skeleton gesture recognition, notably without relying on pre-existing data.
3.1.3 Continuous Gesture Recognition
However, in real-world environments, each video clip doesn’t contain just one gesture and there can be multiple gestures in the continuous video stream. Therefore, we also need to figure out when each gesture starts and ends simultaneously. Building on the foundation established by research in isolated gesture recognition, early endeavors in continuous gesture recognition aimed to build multi-stage pipelines to separate gesture sequences into distinct isolated gesture features for classification. Notable methods include employing a two-stream RNN [109] or 3D CNN [110] for classification, and leveraging 3D CNNs to derive spatial feature maps subsequently integrated with LSTM for analysis [111]. Bhatnagar et al. [112] further applied the dynamic neural network to select features for 3D CNN to recognize continuous gestures. In addition, extracting additional features from depth maps within RGB-D videos, as well as analyzing human poses or facial expressions, can also significantly enhance the performance of continuous gesture recognition. In this area, Zhou et al. [113] introduced a spatial-temporal multi-cue network, a sophisticated approach that amalgamates various features including human poses, gestures, and facial expressions. Besides, Liu et al. [114] developed a two-stream CNN framework that utilizes RGB-D video inputs to accurately detect and classify continuous gestures, showing the potential of combining depth maps for continuous gesture recognition.
3.1.4 Discussion
3.2 Gaze and Attention Analysis
3.2.1 Social Background
3.2.2 Gaze Target Detection
Several works have addressed the task of gaze target detection, which aims to localize the target of a person’s visual attention in an image containing the person and the scene. Early works in this direction combined head-tracking with estimation of head orientation to identify when a person is looking at other people in meetings [120, 121]. Building upon a larger body of work in visual saliency prediction [122], Recasens et al. [115] published a seminal work introducing the GazeFollow benchmark. They proposed a 2-stream deep neural network with separate branches to estimate gaze direction and scene saliency, combining their outputs to produce a gaze heatmap over the scene for a given person. Several works have improved upon this 2-stream paradigm [123, 124, 125, 126, 127, 128], including adapting to 360-degree images [129]. Notably, Chong et al. [116] introduced the VideoAttentionTarget dataset, which provides gaze target annotations for video clips. They proposed a 2-stream model with an Convolutional LSTM [130] to integrate temporal information. Many recent works have demonstrated performance gains by augmenting 2-stream models with pretrained models for estimating additional signals like scene depth [131, 132, 117], body pose [132, 133], head pose [131, 134], and eye detection [131]. A set of recent works has followed a detection framing of the problem, where a set of head bounding boxes and paired gaze targets are detected end-to-end via a DETR-based [135] Transformer network [136, 137, 138].
Other works predict gaze targets from the egocentric point of view. These methods use video from a headworn camera to predict the gaze of the camera wearer. This problem also differs from eye-tracking, which uses close-up cameras capturing the eyes to determine gaze direction; rather, gaze is localized from the egocentric video stream, which captures the scene. Fathi et al. [139] introduced the GTEA and GTEA+ datasets of daily activities with ground truth eye-tracking data, and proposed a probabilistic model for inferring gaze conditioned on action. Further methods predict gaze based on saliency, hand, and egomotion features [140, 141]. Huang et al. and Li et al. proposed deep learning methods for jointly modeling visual attention and activity [142, 143, 144]. Lai et al. [145, 146] proposed a Transformer-based model and further evaluated on the Ego4D [147] eye-tracking subset, which focuses on multi-party social interactions. A related task is anticipating egocentric gaze in future frames [148, 149, 150].
3.2.3 Multi-party Gaze Analysis
A lesser studied direction of gaze analysis has focused on recognizing multi-person social gaze dynamics. In contrast to gaze target detection, tasks in this direction involve reasoning at a higher-level about people and their respective gaze targets in relation to others. One such problem is recognizing shared attention, or when 2 people in a scene are looking at the same target [151, 152, 153, 154]. Park et al. defined the related task of social saliency prediction, which involves identifying potential targets of shared attention in scenes with multiple people [155, 156, 157]. They propose using multiple first-person videos to reconstruct the scene and learn a social saliency model that is applied downstream to third-person images. Prior works have also addressed the task of determining when people look at eachother, referred to as mutual gaze detection or “Looking at Eachother” (LAEO) [158, 159, 160, 161, 162, 163, 164]. Muller et al. propose using gaze and speaking behavior together to identify eye-contact in multi-party interactions [165]. In 2019, Fan et al. [166] introduced the VACATION benchmark with various types of gaze behaviors such as shared attention, mutual gaze and gaze following. They proposed a spatio-temporal graph approach for capturing gaze dynamics between people over time. Further works have developed end-to-end Transformer-based architectures for identifying these gaze behaviors [167, 154].
Researchers have also addressed identifying social gaze dynamics in egocentric videos. Works have studied the task of detecting when a person visible in the egocentric video stream is looking at the camera wearer. Methods have addressed eye-contact detection [168, 169, 170, 171, 172, 173, 174], with several focusing on the use-case of assessing child behavior in the clinical setting [168, 170, 174]. Similarly, the Ego4D dataset [147] includes the “Looking at Me” task, and the related audiovisual “Talking to Me” task as part of their social benchmark. Paralleling the task of identifying shared attention in third-person videos, researchers have explored identifying objects of shared attention across multiple egocentric videos by identifying similarities in views [175, 176, 177]. Fathi et al. [178] addressed identifying social interaction types via inferring the visual attention of both the camera-wearer and people in their view, including dialogue, monologue, and discussion .
3.2.4 Discussion
3.3 Facial Expression Analysis
3.3.1 Social Background
Facial expressions serve as a vital channel for humans to communicate their emotions, playing a key role in the fabric of social interactions. Facial expressions are more than just transient displays of emotion; they are the silent language through which individuals communicate a wide range of feelings, from joy and surprise to sadness and anger, often without uttering a single word. This section explores advancements in facial expression analysis and underscores its importance in cultivating relationships that are sensitive to emotional and social nuances.
In the field of computer vision, numerous facial expression recognition approaches have been developed to interpret expressions from human faces. The origins of it can be traced back to the 20th century, where Ekman and Friesen [182] identified six core emotions—anger, disgust, fear, happiness, sadness, and surprise—proposing that these basic emotions are universally recognized across different cultures. Subsequently, contempt was acknowledged as a seventh fundamental emotion, further broadening our comprehension of human emotional expressions [183]. Although there are various models for categorizing emotions, such as the facial action coding system  [184] and the continuous model based on affect dimensions [185], the categorical model is the most widely adopted approach.
3.3.2 Static Facial Expression Recognition
In earlier times, traditional approaches [186, 187, 188, 189] relied on manually crafted features [190, 191, 192]. However, with the accumulation of sufficient training data since 2013, methods based on deep learning have come into the spotlight, marking a significant shift in the field. In their pioneering work, Savchenko et. al [193] demonstrated the potential of CNNs in facial expression recognition. Following this path, Zhao et. al [194] introduced EfficientFace, where a local-feature extractor and a channel-spatial modulator have been developed to capture local facial characteristics and global-salient features, enhancing the model’s resilience to occlusions and variations in pose. However, a notable limitation of convolution-based approaches is their inability to capture the global context of images, owing to the confined scope of the convolutional local receptive field. Addressing this gap, Xue et. al [195] took inspiration from the vision transformer [196], crafting the first transformer-based network. This network was specifically engineered to grasp long-range dependencies within images, marking a significant stride in this field. Building on this advancement, Kim et. al [197] refined the vision transformer (ViT), enhancing its capability to amalgamate both global and local image features. Following prior research, POSTER [198] tackles the main challenges of facial expression recognition, namely inter-class similarity, intra-class discrepancy, and sensitivity to scale by cleverly integrating facial landmarks with image features using a dual-stream pyramidal cross-fusion transformer architecture. Nonetheless, the computational expense incurred by the network’s architecture impacts the efficiency of facial expression recognition. To mitigate this problem, POSTER++ [199] has been introduced. This approach reconstructs the network structure by streamlining the cross-fusion and multi-scale extraction processes, achieving state-of-the-art performance.
3.3.3 Dynamic Facial Expression Recognition
Advancements in static facial expression recognition (SFER) are now being complemented by growing interest in dynamic facial expression recognition (DFER), which adds the complexity of analyzing temporal alongside spatial features in images. Several techniques[200, 201] leverage CNNs to extract spatial features from frames, then apply RNNs for temporal analysis. 3D CNNs are also used as an alternative. Recently, transformers have emerged as effective tools for extracting both spatial and temporal data, with Zha et al.’s [202] dynamic facial expression recognition transformer (Former-DFER) incorporating a convolutional spatial transformer and a temporal transformer. Ma et al. [203] introduced the spatial-temporal Transformer (STT) for capturing frame-specific features and their contextual relationships. Additionally, Wang et al. [204] developed the Dual Path multi-excitation Collaborative Network (DPCNet) to learn facial expression features from key frames.
3.3.4 Discussion
Facial expressions are a key channel for conveying emotions and social signals in human interactions. However, current facial recognition systems often struggle with biases related to gender, race, and cultural background [205, 206, 207, 208]. These biases can lead to the misinterpretation or misuse of facial expression analysis in social contexts, perpetuating stereotypes and discrimination against underrepresented groups. While various approaches have been proposed to mitigate these biases, effectively addressing them remains a significant challenge [209, 210, 211, 212, 213]. Future research should develop more inclusive and unbiased facial expression recognition systems. In addition, privacy is an important issue in systems that handle facial data. The use of these systems without proper regulation can lead to the violation of privacy rights and the misuse of their personal information [214]. Therefore, it is also important to develop AI systems that consider privacy protection and ensure the responsible use of facial data.
@@ 3.3.4 Discussion
3.4 Datasets with Non-verbal Cues
Table II shows a comprehensive overview, comparing the attributes of datasets with non-verbal cues. Further details for each dataset are included in the supplementary material.
4 Understanding Multimodal Cues
4.1 Multimodal Emotion Analysis
4.1.1 Social Background
Human communication is inherently multimodal, conveying emotions through verbal cues like spoken utterances, as well as non-verbal behaviors including gestures, gaze, and facial expressions. Therefore, it is required to understand multimodal cues for recognizing emotional expressions effectively. For instance, it is difficult to detect sarcasm based purely on utterance texts [232]. However, humans can excel at this task by recognizing tone changes and interpreting body language like shrugging shoulders or curling lips. Such multimodal nature adds complexity to emotion analysis [233], and multimodal emotion recognition aims to achieve a more comprehensive and nuanced understanding of human emotional states by leveraging multimodal cues.
4.1.2 Multimodal Emotion Recognition
Unlike early stages of emotion analysis research, which predominantly focused on single modalities like speech [234, 235], textual content [236, 237] and facial expressions [238, 239], the development of multimodal emotion recognition underscores the advantages of jointly considering various channels of expression simultaneously to achieve accurate sentiment recognition[240, 241, 242]. The emotion recognition in conversation task, contextual information from different modalities becomes particularly indispensable for decoding emotional expression and interpretation [231]. Poria et al. [243] propose an LSTM-based model that fuses textual, audio, and visual features to capture interaction history context. Hazarika et al. [45] introduce a model that uses two speaker-dependent GRUs as memory to model utterance context. As an extension to CMN, [244] adds an extra global GRU to track the inter-speaker dependency that exists in the entire conversation. However, neither methods can model multiparty conversations involving more than two speakers because the number of memory network blocks is fixed. Besides, neither of them can distinguish individuals according to given utterances.
[…]
4.1.3 Discussion
4.2 Conversation Dynamics Analysis
Figure 10: Examples of conversation dynamics analysis in multi-party settings from [248]. Through the consideration of multimodal cues, it is possible to understand the dynamics of the conversation such as who is speaking to whom and who is referred in multi-party environments well.
4.2.1 Social Background
Conversation dynamics analysis aims to model the fine-grained interactions among multiple people engaged in a conversation, which is accompanied by modeling the interplay of verbal and non-verbal cues. It involves modeling various aspects of dynamics such as identifying who is speaking to whom, inferring who is the target of attention, and capturing the referential relationships between speakers and listeners. Effective communication requires interpreting these subtle dynamics, which is essential for machines to seamlessly understand multi-party interactions.
4.2.2 Conversation Dynamics Understanding
To explicitly study gaze behaviors with multimodal cues, Hou et al. [249] introduced a multimodal gaze following model for conversational scenarios. In the egocentric analysis context, Jiang et al. [250] involved the multi-channel audio stream to localize the active speaker in all possible directions on a sphere. Ryan et al. [251] proposed the new task of selective auditory attention localization to detect who the camera wearer is listening to using egocentric videos and multi-channel audios. They introduced a transformer-based framework to model the relationships between audiovisual features in different spatial regions. Jia et al. [252] introduced the concept of an audio-visual conversational graph to represent the conversational behaviors (speaking and listening) of both the camera wearer and all social partners. Regarding the language-visual modeling, Lee et al. [248] tackled the problem of identifying referents in multiparty conversations, curating new social tasks: speaking target identification, pronoun coreference resolution, and mentioned player prediction. They proposed a model that established dense language-visual alignments by synchronizing player visual features with their corresponding utterances.
Some other works in this area concentrates on modeling the poses and spatial relation of participants in a conversational group, including inferring head, body orientation and facing formations, as well as detecting the independent conversational groups in cocktail party  [253, 254, 255, 256]. Ricci et al. [253] proposed a coupled learning framework to jointly estimate head pose, body pose, and F-formations from surveillance videos while handling occlusions. Alameda-Pineda et al. [254] introduced SALSA, which records a social event using cameras and wearable sensors, discussing challenges in crowded interactions. Tan et al. [255] presented an LSTM-based model that predicts continuous pairwise affinities to identify conversational groups, leveraging spatial context and temporal dynamics in positional and orientation cues. Joo et al. [257] introduced a framework to predict social signals such as speaking status, social formation, and body gestures in triadic social interactions based on 3D motion capture data. To this end, they leveraged various types of social cues such as body motions, face expressions, positions, and voice signals.
4.2.3 Discussion
4.3 Social Situation Analysis
4.3.1 Social Background
4.3.2 Social Relationship Reasoning
4.3.3 Social Question Answering
Social-IQ [258] is a Q&A benchmark for understanding social situations. It features various in-the-wild social videos, each with questions and multiple-choice answers. The questions address: (1) causes and intentions behind a social situation, (2) how and why humans act in a certain manner, (3) mental state, trait, attitude, and (4) memory for referencing. The task is to choose the correct answer based on social understanding. Sartzetaki et at. [269] developed a model which embeds the visual, textual context and audio inputs with three separate encoders. The question is encoded by a control unit and then injected into the embedding of each modality to learn a question-related representation. Wilf et al. [270] introduced the graphic model into social intelligence answering. They encode the multimodal input of each speaking turn into a few nodes in a graph and use a factorization node to merge the information. Xie et al. [271] proposed to leverage emotional characteristics as common sense to select positive and hard negative samples for contrastive learning. They finetune a text encoder on emotion recognition dataset to measure the distance of each answer and the question, which determines the positive and negative pairs for training. Natu et al. [272] used language models to generate some facts for each video based on the transcript. These facts are integrated into the multimodal model as global common sense by straightforward concatenation or treating it as an additional modality. Zhu et al. [273] further annotated personality labels in TVQA [274]. Then they proposed two problems – predicting (1) the MBTI personality based on the video and (2) the response of a specific character given their personality.
4.3.4 Other Social Situation Analyses
In addition to the relationship reasoning and question answering, some work models the social situation analysis in other forms. Hyun et al. [275] collected a multimodal dataset, named SMILE, which encompasses video clips including multi-party scenarios. Each video clip is annotated with a reason explaining why the person is laughing. Different than question-answering problem that requires selecting the correct answer from candidates, they generate the laughter reason in a free-form language and measure the performance by calculating the similarity of the generated reason and the annotation. Lai et al. [11] established a multimodal video dataset of people playing the werewolf social deduction game. They propose two problems: (1) predicting which persuasion strategies are used in each utterance and (2) predicting the voting of each player.
4.3.5 Discussion
Social situation analysis aims to provide a holistic understanding of the context, norms, and relationships in social interactions. While question answering has been a primary benchmark for evaluating social situation understanding, it has limitations in capturing the complexity of real-world social situations. There also may be some biases between questions and answers [276]. While alternative approaches like modeling social relationships and utterance strategies provide complementary insights, they still face challenges in addressing potential data bias and incorporating common sense and cultural contexts. The development of more comprehensive and unbiased benchmarks, such as those that incorporate free-form understanding, can be helpful for advancing the field. Moreover, the integration of cultural knowledge and social norms into social situation analysis models can help capture the nuances and variations in social behaviors across different contexts.